If the average speed of a light plane had been 40km/h, the plane would have taken one hour longer to fly 1200 km. What was the speed of the plane?
Answer by stanbon(75887)

If the average speed of a light plane had been 40km/h, the plane would have taken one hour longer to fly 1200 km. What was the speed of the plane?
--------------------
"If" DATA:
rate = 40 km/h ; distance = 1200 km ; time = 1200/40 = 30 hrs.
-------------------------------
Actual flight DATA:
rate = x km/h ; distance = 1200 km ; time = 1200/x hrs
-------------------------------
"If" time - Actual time = 1 hr
30 - 1200/x = 1
1200/x = 29
x = 1200/29
x = 41.38 km/hr (Actual speed of the plane)
=====================
Cheers,
Stan H.
=================